OpenStudy (happykiddo):

l s-5 l= 8 l t+2 l= 10 Which of the following is a possible value of s + t? a.) -4 b.) -3 c.) 1 d.) 13 e.) 25

OpenStudy (sparrow2):

it's c so when t=-12 and s=13

OpenStudy (loser66):

Question: 1s-5... what?? = 8

OpenStudy (anonymous):

Remove absolute and put -+ answer on the right side. You find t and s values :)

OpenStudy (loser66):

The give equations are about s and l, while question is about s and t. I don't get it.

OpenStudy (perl):

l s-5 l= 8 gives us two equations: s-5 = 8 s-5 = -8 or s = 13 s = -3 l t+2 l= 10 gives us two equations t+2= 10 t+2 = -10 or t = 8 t = -12 s + t = 13 + 8, 13 + (-12), -3 + 8, -3 + (-12)

OpenStudy (happykiddo):

thanks perl

OpenStudy (loser66):

oh, absolute value!!! OMG

OpenStudy (triciaal):

@Loser66 it's not the letter l

OpenStudy (triciaal):

yeah

OpenStudy (loser66):

hahaha... I am a big idiot>!!!

OpenStudy (triciaal):

no it looks confusing